1) For the following reaction, which of the following is a conjugate acid-base pair? HC20(aq)+ H20) H:0 (aq) C0,(aq) a. HC20 and H20 b. H20 and C202 c. HC20 and H30 d. HCzOr and C20 2) The pH of a solution of hydrochloric acid is 1.5. What is the concentration of the acid? a. 0.22 mol/dm b. 0.32 mol/dm3 c. 3.2x 102 mol/dm d. 2.2x 10-3 mol/dm 3) Which of the following diagrams represents the acid with the highest pH? HINT: arrange the fo boxes from lowest to highest concentration of acid. b. 4) What is the molarity of a solution that contains 22 grams of NasPO4 (Mw 163.94) dissolved in water with the total volume of 300 mL a. 0.54 AM b. 0.45 M c. 0.89 M d. 0.19 AM e. 0.36 M 5) What is the conjugate acid of NaaHAso, a. H30+ b. NasAs0 c. NaH2As0 d. H20 e. HaAsO
Answer #1

1.

(d) HC2O4- and C2O42- are conjugate acid-base pair.

2.

HCl strong acid. [H+]= 10-1.5

Concentration = 10-1.5 M= 3.16 x 10-2 M

The answer is (c)

3.The diagram is not clear.

4. moles = 22/163.94= 0.1342

Molarity= 0.1342x 1000/300 =0.447 M =0.45 M

The answer is (b).

5. The conjugate acid is NaH2AsO4 (c).

The answer is (c).